OpenStudy (anonymous):

Finish writing the equation of the line ____, given the points (-12, 0) and (4, 4). Type only the part that is missing in the equation.

OpenStudy (anonymous):

slope = 4 / 16 = 1/4 using y-y1 = m(x-x1) y-4 = 1/4 (x - 4) 4y - 16 = x - 4 4y - x = 12